1. Two billiard balls (a striped and a solid) coll
Answer #1

Momentum = 20-6 =14 kgm/s (in oppopsite direction of initial velocity)

potential energy = mgh = 10*9.81*4 =392.5 Joules

Kinetic energy = mv^2/2 = 2420 Joules